OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that is parallel to the given line and that passes through the given point. y=2x-4; (1, -1) Some help please?!

OpenStudy (anonymous):

Parallel lines have the same slope; therefore, m = 2 Point-slope form of a line: \[y-y1=m(x-x1)\] \[y+1=2(x-1)\] \[y=2x-1\]

OpenStudy (anonymous):

y=2x-3, sorry

OpenStudy (anonymous):

If two lines are parallel, then they have the same slope. The slope of your lines is 2. So your line should look like 2x+m. 2+m=-1 => m=-3. So your equation is y=2x-3.